### v3

@@ 8 Conclusion
${\bf E}\bigg{[}\bigg{|}\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}]){\bf E}[Z_{j}]A_{ij}(1)\bigg{|}\bigg{]}\leq{\bf E}\bigg{[}\bigg{(}\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}])V_{i}\bigg{)}^{2}\bigg{]}^{1/2}\\ =\bigg{(}\frac{k_{n}^{2}}{n^{4}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\Lambda_{i}}\text{Cov}(Z_{i},Z_{j})V_{i}V_{j}+\frac{k_{n}^{2}}{n^{4}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\not\in\Lambda_{i}}\text{Cov}(Z_{i},Z_{j})V_{i}V_{j}\bigg{)}^{1/2}.$ (SA.6.31)
@@
(SA.6.31)$\precsim\bigg{(}\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\Lambda_{i}}\lvert\text{Cov}(Z_{i},Z_{j})\rvert+\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\not\in\Lambda_{i}}\lvert\text{Cov}(Z_{i},Z_{j})\rvert\bigg{)}^{1/2}.$
@@
By Assumption 5(a), $\Lambda_{i}\subseteq\mathcal{N}(i,2(r_{n}+U_{n}))\subseteq\mathcal{N}(i,3U_{n})$, and $\Lambda_{i}\backslash C_{c(i)}\subseteq\mathcal{N}(i,3U_{n})\backslash\mathcal{N}(i,L_{n})$. By Assumption 5(b), there exists a non-negative sequence $\alpha_{n}\precsim 1$ such that $U_{n}-L_{n}=\alpha_{n}(n/(k_{n}\xi_{n}))^{1/d}$. Then by Assumption 2(b), following (SA.6.15),
$\max_{i}\lvert\mathcal{N}(i,3U_{n})\backslash\mathcal{N}(i,L_{n})\rvert\precsim\alpha_{n}\xi_{n}(n/(k_{n}\xi_{n}))^{(d-1)/d}.$
@@
$\frac{k_{n}^{2}}{n^{4}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}\sum_{k\in\mathcal{N}_{n}}\sum_{\ell\in\mathcal{N}_{n}}\text{Cov}({\bf E}[Z_{i}\mid\mathcal{F}_{i}]{\bf E}[Z_{j}\mid\mathcal{F}_{j}],{\bf E}[Z_{k}\mid\mathcal{F}_{k}]{\bf E}[Z_{\ell}\mid\mathcal{F}_{\ell}])A_{ij}(1)A_{k\ell}(1).$ (SA.6.32)
@@
(SA.6.32)$\precsim\frac{k_{n}^{2}}{n^{4}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\Lambda_{i}}\lvert\Xi_{ij}\rvert$ (SA.6.33)
@@
(SA.6.33)$\precsim\frac{k_{n}^{2}}{n^{4}}\cdot n\cdot\frac{n}{k_{n}}\cdot\frac{n^{2}}{k_{n}^{2}}\precsim k_{n}^{-1}\prec 1.$
Step 5. Recall the definition of $\tilde{\sigma}_{n}^{2}$ from (SA.6.28). We next prove that
@@
Case 1. $\min\{\lvert C_{1}\rvert,\lvert C_{2}\rvert\}\precsim n/k_{n}$. Then the argument largely proceeds as in the proof of Theorem 3 of Cao et al. (2022) but with additional accounting for $k_{n},\xi_{n}$ diverging. Since clusters partition $\mathcal{N}_{n}$, for every $n$ there must exist some cluster $C_{3}$ of size at least $n/k_{n}$. Let $i_{3}$ denote its medoid and $R_{3}$ be the largest distance between $i_{3}$ and an element in $C_{3}$. By Assumption 2(a),
$\frac{n}{k_{n}}\leq\lvert C_{3}\rvert\leq\lvert\mathcal{N}(i_{3},R_{3})\rvert\leq C\xi_{n}R_{3}^{d}\quad\Longrightarrow\quad R_{3}\succsim(n/(k_{n}\xi_{n}))^{1/d}.$ (SA.6.34)
@@
Consider a hypothetical update in step 2 of Algorithm SA.1.1 that replaces $i_{2}$ with $i_{3}^{\prime}$. Then all units in $\mathcal{N}(i_{3}^{\prime},R_{3}/8)$ are optimally reassigned to the cluster with medoid $i_{3}^{\prime}$ because they are all by construction at least distance $3/8\cdot R_{3}$ away from any other medoid. This reassignment reduces the total cost by at least $\lvert\mathcal{N}(i_{3}^{\prime},R_{3}/8)\rvert R_{3}/4\succsim n/k_{n}\cdot(n/(k_{n}\xi_{n}))^{1/d}$ by Assumption 2(a) and (SA.6.34). On the other hand, in the worst case, all other units in $C_{2}$ are reassigned to medoid $i_{1}$, which increases the total cost by at most $\lvert C_{2}\rvert\ell_{n}(n/(k_{n}\xi_{n}))^{1/d}\prec n/k_{n}\cdot(n/(k_{n}\xi_{n}))^{1/d}$. Hence, the update is overall cost-reducing for $n$ sufficiently large, which contradicts for such $n$ the supposition that the clusters are the output of Algorithm SA.1.1, in particular violating step 2.
Case 2. $\min\{\lvert C_{1}\rvert,\lvert C_{2}\rvert\}\succ n/k_{n}$. Without loss of generality, suppose $\lvert C_{2}\rvert\succ n/k_{n}$. Then the argument proceeds similarly to case 1 but using $C_{2}$ in place of $C_{3}$. In particular, let $R_{2}$ be the largest distance between $i_{2}$ and an element of $C_{2}$. By an argument similar to (SA.6.34), $R_{2}\succ(n/(k_{n}\xi_{n}))^{1/d}$.
@@
Recall from Assumption 5 the definition of cluster centroids, $U_{n}$, and $L_{n}$. For any $i\in\mathcal{N}_{n}$, all clusters intersecting $\mathcal{N}(i,r_{n})$ must be subsets of $\mathcal{N}(i,r_{n}+2U_{n})$. By Assumption 2(a), Assumption 5, and (2), there exists $\alpha_{n}\sim 1$ such that $\mathcal{N}(i,r_{n}+2U_{n})\subseteq\mathcal{N}(i,\alpha_{n}L_{n})$ for any $i\in\mathcal{N}_{n}$. Then Assumption 5 implies that $\phi_{i}$ is at most the number of $L_{n}$-balls that fill $\mathcal{N}(i,\alpha_{n}L_{n})$ without intersecting. We therefore seek to bound the $2L_{n}$-packing number (Wainwright, 2019, Definition 5.4) of $\mathcal{N}(i,\alpha_{n}L_{n})$. By Lemma 5.5 of Wainwright (2019), this is at most the $L_{n}$-covering number of the ball (Wainwright, 2019, Definition 5.1), and we denote this number by $N_{n}(i)$.
To bound $N_{n}(i)$, we follow the proof of Lemma 5.7(b) in Wainwright (2019). Construct a maximal $2L_{n}$-packing of $\mathcal{N}(i,\alpha_{n}L_{n})$ with cardinality $M$ and centroids $\{\theta_{m}\}_{m=1}^{M}$. This is also an $L_{n}$-covering of $\mathcal{N}(i,\alpha_{n}L_{n})$, so $N_{n}(i)\leq M$. The balls of the packing $\{\mathcal{N}(\theta_{m},L_{n}/2)\}_{m=1}^{M}$ are disjoint and contained in $\mathcal{N}(i,\alpha_{n}L_{n}+L_{n}/2)$, so
$\sum_{m=1}^{M}\lvert\mathcal{N}(\theta_{m},L_{n}/2)\rvert\leq\lvert\mathcal{N}(i,L_{n}(\alpha_{n}+0.5))\rvert.$
By Assumption 2(a), there exists $C>0$ not dependent on $i$ such that
$\sum_{m=1}^{M}\lvert\mathcal{N}(\theta_{m},L_{n}/2)\rvert\geq\frac{M}{C}\xi_{n}\left(\frac{L_{n}}{2}\right)^{d}\quad\text{and}$
@@
Hence, $M\leq C^{2}(1+2\alpha_{n})^{d}\precsim 1$, which does not depend on $i$, so $\max_{i}N_{n}(i)\precsim 1$, which proves the first claim of the lemma.
@@
By Lemma SA.5.3, $p_{ti}$ is bounded away from zero uniformly over $i$ and by Assumptions 2(a) and 5, $\max_{j}\lvert C_{j}\rvert\precsim n/k_{n}$, so the right-hand side is at most of order

### v4

@@ 8 Conclusion
${\bf E}\bigg{[}\bigg{|}\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}]){\bf E}[Z_{j}]A_{ij}(1)\bigg{|}\bigg{]}\leq{\bf E}\bigg{[}\bigg{(}\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}])V_{i}\bigg{)}^{2}\bigg{]}^{1/2}\\ =\bigg{(}\frac{k_{n}^{2}}{n^{4}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\Lambda_{i}}\text{Cov}(Z_{i},Z_{j})V_{i}V_{j}+\frac{k_{n}^{2}}{n^{4}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\not\in\Lambda_{i}}\text{Cov}(Z_{i},Z_{j})V_{i}V_{j}\bigg{)}^{1/2}.$ (SA.6.30)
@@
(SA.6.30)$\precsim\bigg{(}\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\Lambda_{i}}\lvert\text{Cov}(Z_{i},Z_{j})\rvert+\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\not\in\Lambda_{i}}\lvert\text{Cov}(Z_{i},Z_{j})\rvert\bigg{)}^{1/2}.$
@@
By Assumption 5(a), $\Lambda_{i}\subseteq\mathcal{N}(i,2(r_{n}+U_{n}))\subseteq\mathcal{N}(i,3U_{n})$, and $\Lambda_{i}\backslash C_{c(i)}\subseteq\mathcal{N}(i,3U_{n})\backslash\mathcal{N}(i,L_{n})$. By Assumption 5(b), there exists a non-negative sequence $\alpha_{n}\precsim 1$ such that $3U_{n}-L_{n}=\alpha_{n}(n/(k_{n}\xi_{n}))^{1/d}$. Then by Assumption 2(b), following (SA.6.15),
$\max_{i}\lvert\mathcal{N}(i,3U_{n})\backslash\mathcal{N}(i,L_{n})\rvert\precsim\xi_{n}(n/(k_{n}\xi_{n}))^{(d-1)/d}.$
@@
$\frac{k_{n}^{2}}{n^{4}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}\sum_{k\in\mathcal{N}_{n}}\sum_{\ell\in\mathcal{N}_{n}}\text{Cov}({\bf E}[Z_{i}\mid\mathcal{F}_{i}]{\bf E}[Z_{j}\mid\mathcal{F}_{j}],{\bf E}[Z_{k}\mid\mathcal{F}_{k}]{\bf E}[Z_{\ell}\mid\mathcal{F}_{\ell}])A_{ij}(1)A_{k\ell}(1).$ (SA.6.31)
@@
(SA.6.31)$\precsim\frac{k_{n}^{2}}{n^{4}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\Lambda_{i}}\lvert\Xi_{ij}\rvert$ (SA.6.32)
@@
(SA.6.32)$\precsim\frac{k_{n}^{2}}{n^{4}}\cdot n\cdot\frac{n}{k_{n}}\cdot\frac{n^{2}}{k_{n}^{2}}\precsim k_{n}^{-1}\prec 1.$
Step 5. Recall the definition of $\tilde{\sigma}_{n}^{2}$ from (SA.6.27). We next prove that
@@
Case 1. $\min\{\lvert C_{1}\rvert,\lvert C_{2}\rvert\}\precsim n/k_{n}$. Then the argument largely proceeds as in the proof of Theorem 3 of Cao et al. (2024) but with additional adjustments to account for the divergence of $k_{n},\xi_{n}$. Since clusters partition $\mathcal{N}_{n}$, for every $n$ there must exist some cluster $C_{3}$ of size at least $n/k_{n}$. Let $i_{3}$ denote its medoid and $R_{3}$ be the largest distance between $i_{3}$ and an element in $C_{3}$. By Assumption 2(a),
$\frac{n}{k_{n}}\leq\lvert C_{3}\rvert\leq\lvert\mathcal{N}(i_{3},R_{3})\rvert\leq C\xi_{n}R_{3}^{d}\quad\Longrightarrow\quad R_{3}\succsim(n/(k_{n}\xi_{n}))^{1/d}.$ (SA.6.33)
@@
Consider a hypothetical update in step 2 of Algorithm SA.1.1 that replaces $i_{2}$ with $i_{3}^{\prime}$. Then all units in $\mathcal{N}(i_{3}^{\prime},R_{3}/8)$ are optimally reassigned to the cluster with medoid $i_{3}^{\prime}$ because they are all by construction at least distance $3/8\cdot R_{3}$ away from any other medoid. This reassignment reduces the total cost by at least $\lvert\mathcal{N}(i_{3}^{\prime},R_{3}/8)\rvert R_{3}/4\succsim n/k_{n}\cdot(n/(k_{n}\xi_{n}))^{1/d}$ by Assumption 2(a) and (SA.6.33). On the other hand, in the worst case, all other units in $C_{2}$ are reassigned to medoid $i_{1}$, which increases the total cost by at most $\lvert C_{2}\rvert\ell_{n}(n/(k_{n}\xi_{n}))^{1/d}\prec n/k_{n}\cdot(n/(k_{n}\xi_{n}))^{1/d}$. Hence, the update is overall cost-reducing for $n$ sufficiently large, which contradicts for such $n$ the supposition that the clusters are the output of Algorithm SA.1.1, in particular violating step 2.
Case 2. $\min\{\lvert C_{1}\rvert,\lvert C_{2}\rvert\}\succ n/k_{n}$. Without loss of generality, suppose $\lvert C_{2}\rvert\succ n/k_{n}$. Then the argument proceeds similarly to case 1 but using $C_{2}$ in place of $C_{3}$. In particular, let $R_{2}$ be the largest distance between $i_{2}$ and an element of $C_{2}$. By an argument similar to (SA.6.33), $R_{2}\succ(n/(k_{n}\xi_{n}))^{1/d}$.
@@
Recall from Assumption 5 the definition of cluster centroids, $U_{n}$, and $L_{n}$. For any $i\in\mathcal{N}_{n}$, all clusters intersecting $\mathcal{N}(i,r_{n})$ must be subsets of $\mathcal{N}(i,r_{n}+2U_{n})$ by the assumption. Together with (2), we have that there exists $\alpha_{n}\sim 1$ such that $\mathcal{N}(i,r_{n}+2U_{n})\subseteq\mathcal{N}(i,\alpha_{n}L_{n})$ for any $n$ and $i\in\mathcal{N}_{n}$. Invoking Assumption 5 once more, it follows that $\phi_{i}$ is at most the number of $L_{n}$-balls that fill $\mathcal{N}(i,\alpha_{n}L_{n})$ without intersecting.
We seek to bound the $2L_{n}$-packing number (Wainwright, 2019, Definition 5.4) of $\mathcal{N}(i,\alpha_{n}L_{n})$. By Lemma 5.5 of Wainwright (2019), this is at most the $L_{n}$-covering number of the ball (Wainwright, 2019, Definition 5.1), and we denote this number by $N_{n}(i)$. We bound this following the proof of Lemma 5.7(b) in Wainwright (2019).
Construct a maximal $L_{n}/2$-packing of $\mathcal{N}(i,\alpha_{n}L_{n})$ with cardinality $M_{i}$ and centroids $\{\theta_{m}\}_{m=1}^{M_{i}}$. This is also an $L_{n}$-covering of $\mathcal{N}(i,\alpha_{n}L_{n})$, so $N_{n}(i)\leq M_{i}$. The balls of the packing $\{\mathcal{N}(\theta_{m},L_{n}/2)\}_{m=1}^{M_{i}}$ are disjoint and contained in $\mathcal{N}(i,\alpha_{n}L_{n}+L_{n}/2)$, so
$\sum_{m=1}^{M_{i}}\lvert\mathcal{N}(\theta_{m},L_{n}/2)\rvert\leq\lvert\mathcal{N}(i,L_{n}(\alpha_{n}+0.5))\rvert.$
By Assumption 2(a), there exists $C>0$ independent of $i$ and $n$ such that
$\sum_{m=1}^{M_{i}}\lvert\mathcal{N}(\theta_{m},L_{n}/2)\rvert\geq\frac{M_{i}}{C}\xi_{n}\left(\frac{L_{n}}{2}\right)^{d}\quad\text{and}$
@@
Hence, $M_{i}\leq C^{2}(1+2\alpha_{n})^{d}\precsim 1$, so $\max_{i}N_{n}(i)\precsim 1$, which proves the first claim of the lemma.
@@
By Lemma SA.5.3, $\min_{i}p_{ti}\succsim 1$, and by Assumptions 2(a) and 5, $\max_{j}\lvert C_{j}\rvert\precsim n/k_{n}$, so the right-hand side is at most of order